In two years time James will be half the age of his sister, who is now 12. How old is James now?

Answer by macston(5194)

J=age of James; S=age of sister=12
J+2yrs=1/2(S+2yrs)
J+2yrs=1/2(12yrs+2yrs)
J+2yrs=1/2(14yrs)
J+2yrs=7yrs
J=7yrs-2yrs=5 yrs ANSWER: James is now 5 years old.